# Tuning the band topology and topological Hall effect in skyrmion crystals via the spin-orbit coupling

### v1

@@ Abstract
The topological Hall effect (THE) is the result of spin-asymmetric deflection of charge carriers flowing through a non-collinear spin system. Effective manipulation of the topological Hall conductivity (THC) in skyrmions is currently a vigorous area of research with an eye towards potential spintronics application. Here, we show that the band topology and the THC in a skyrmion crystal can be tuned by changing the strength of the Rashba spin-orbit coupling (SOC), which can be accomplished via a perpendicular electric field. This results in the change of the subband Chern numbers and a transition between ordinary insulator and Chern insulator as the Rashba SOC is varied. For partially filled subbands, the Rashba SOC can tune the THC and reverse its sign, so that the direction of the Hall current is flipped. The critical Rashba strength for this depends on the skyrmion type and the carrier density. We extend our analysis to the cases of Dresselhaus and Weyl SOC as well, and show that they can be directly mapped to the Rashba SOC case and therefore lead to similar results. Our work suggests new avenues for the control of charge transport in skyrmion crystals.
@@ I Introduction
In this work, we study the THE in the skyrmion crystal in the presence of three types of SOC, viz., the Rashba, Dresselhaus, or the Weyl type. The Rashba case is studied in some detail, while we show that the other two cases can be mapped to the Rashba case for the skyrmion types we have considered here. Quite remarkably, we find that not only can the magnitude of the THC be significantly changed by varying the strength of the Rashba SOC via an applied electric field, but also that the sign of the Hall conductivity can be reversed by varying the Rashba SOC around a critical value. The band topology in the skyrmion crystal can also be changed by varying the strength of the SOC term, resulting in a change in the subband Chern numbers, which can lead to a flipping of the direction of the topological edge current.
@@ II Model Hamiltonian and the electron band structure
The first term is the nearest neighbor hopping term, where we have suppressed the spin index, with $c_{i}^{\dagger}\equiv(c_{i\uparrow}^{\dagger},c_{i\downarrow}^{\dagger})$ and $c_{i}\equiv(c_{i\uparrow},c_{i\downarrow})^{T}$, $i$ is the site index on the lattice, $\sigma$ is the spin index, $c_{i\sigma}^{\dagger}/c_{i\sigma}$ are the corresponding electron creation/annihilation operators, $\langle ij\rangle$ indicates distinct pairs of nearest neighbors, and H. c. indicates Hermitian conjugate of the last term only. The second term is the exchange interaction between the site-fixed spins $\bm{S}_{i}$, commonly taken as classical spins, with the skyrmion texture and $\bm{\sigma}$ is the Pauli matrix describing the electron spin. The third term is the lattice version of the Rashba SOC[49, 50], which breaks the $z\rightarrow-z$ symmetry. In this term, $\alpha_{R}$ is the Rashba interaction strength and $\mathbb{r}_{ij}=\mathbb{r}_{j}-\mathbb{r}_{i}$ is the vector distance between the nearest-neighbor sites. It is easily shown that (Supplementary Materials) for the free-electron case (small $k$ limit of the lattice case), it yields the well-known form of the Rashba Hamiltonian ${\cal H}_{R}^{k}=\alpha_{R}^{\prime}(\bm{\sigma}\times\bm{k})\cdot\hat{z}$, with a scaled coefficient $\alpha_{R}^{\prime}=2a^{2}\alpha_{R}$, with $a$ being the lattice constant. The Rashba SOC can arise due to the symmetry breaking due to an overlayer or by applying an external electric field. For the free electrons in the presence of an electric field, the strength of the Rashba term is given by $\alpha_{R}^{\prime}=\hbar^{2}E_{z}/2m_{e}^{2}c^{2}$. However, for the solids, it is enormously enhanced owing to the fact that the relativistic effects originate largely from the nuclear region[51].
The exchange interaction is often taken as $J\rightarrow\infty$, which is reasonable considering that for typical solids $J\sim 2-3$ eV, while the typical hopping is a fraction of an eV. Also, the essential physics discussed here is not changed by taking this limit. However, there are no additional complications if a general $J$ is used. In infinite $J$ limit, the problem reduces to that of the spinless fermions, since the electron state with spin anti-aligned with the lattice spin is forbidden. The spinless fermion operators are described by
@@
where $(\theta,\phi)$ are the polar and azimuthal angles in the spherical polar coordinates and $(r,\alpha)$ is the 2D position of the lattice spin in polar coordinates with the origin coinciding with the skyrmion centre. Here, $\lambda$ is the skyrmion radius, $m$ is the winding number or the vorticity, and $\gamma$ is the helicity of the skyrmion. For a single-turned skyrmion and antiskyrmion, $m$ takes the value of $+1$ and $-1$, respectively. Depending on the helicity, skyrmions can be classified as Néel ($\gamma=0\ {\rm or}\ \pi$) and Bloch types ($\gamma=\pm\pi/2$) as illustrated in Figure 2.
@@
We consider a square lattice of localized spins with a magnetic unit cell of $N\times N$ sites hosting a skyrmion with the spin texture Eq. 4. Thus the skyrmions form a square lattice with lattice constant $Na$, where $a=1$ is the distance between the neighboring spins. We took $N=4$ with the skyrmion radius $\lambda=2a$, so that there are $N^{2}=16$ sites in the magnetic unit cell of the skyrmion crystal. Other skyrmion/antiskyrmion crystals are defined in a similar way.
@@
The SOC modifies the emergent magnetic field of the skyrmion, which we discuss below. In addition, it modifies the magnitude of the effective hopping (see Supplementary Materials for details). Both these effects modify the electronic band structure as well as the topological properties.
@@
Emergent magnetic field. In the lattice models with a magnetic field, hopping integrals between lattice sites acquire a complex phase factor $t\times\exp\big{[}\frac{-ie}{\hbar}\int_{a}^{b}\mathbb{A}.d\mathbb{r}\big{]}$ via the Peierls substitution, where $t$ is the hopping without the magnetic field, $\mathbb{A}$ is the vector potential, $-e<0$ is the charge of the electron, and the integral is taken over the path joining the two sites. To compute the magnetic field, we take an infinitesimal square loop around the point $\mathbb{r}$ and compute the flux through it using the Peierls phase factors accumulated by the electron as it travels around the square loop, following the Hamiltonian Eq. (3). Note that the magnetic field is independent of the lattice type of the skyrmion crystal (square lattice in this paper). After straightforward but tedious algebra, we find that the magnetic field, which is normal to the skyrmion plane, is given by
$B_{SK}^{R}(r)$ $=$ $\frac{\Phi_{0}}{2\pi}\times\big{[}\frac{\pi}{2\lambda r}\sin\frac{\pi r}{\lambda}$ (5)
@@
-- $\frac{\alpha_{R}}{t}(\frac{\pi}{\lambda}\cos{\frac{\pi r}{\lambda}}+\frac{1}{r}\sin\frac{\pi r}{\lambda})\cos\gamma\big{]}$
for $r<\lambda$, and zero for $r>\lambda$, and $\Phi_{0}=h/e$ is the flux quantum. The first term is due to the skyrmion alone and the second term is the modification due to the Rashba SOC. The superscript in $B_{SK}^{R}$ refers to the Rashba SOC case. For the antiskyrmion, the expression is
@@
$B_{AS}^{R}(r,\alpha)$ $=$ $\frac{\Phi_{0}}{2\pi}\times\big{[}-\frac{\pi}{2\lambda r}\sin\frac{\pi r}{\lambda}$ (6)
-- $\frac{\alpha_{R}}{t}(\frac{\pi}{\lambda}\cos{\frac{\pi r}{\lambda}}-\frac{1}{r}\sin\frac{\pi r}{\lambda})\cos({2\alpha}-\gamma)\big{]},$
@@
where the magnetic field is no longer circularly symmetric. It can be shown by direct integration that the total flux passing through a single syrmion/ antiskyrmion is ±$\pm\Phi_{0}$, irrespective of the strength of the Rashba term.
@@
Figure 3: Band structure and THC in the skyrmion crystal with the Néel skyrmion. (a) Band structure without (red lines) and with (blue lines) Rashba SOC. For the latter case, only the lowest four bands are shown. (b) THC as a function of the Fermi energy $\varepsilon_{F}$, indicating prominent Hall plateaus for $\alpha_{R}=0.5t/a$, chosen so as to produce a maximally uniform magnetic field (see Fig. 4).
@@
[…]
@@
[…]
@@
The band structure for the $4\times 4$ lattice with the Rashba strength $\alpha_{R}^{0}$ is shown in Figure 3 (a) (in blue), which is very similar to the bands with a uniform magnetic field (see Supplementary Materials). As discussed in the next Section, we show that the topology of the bands become modified in the presence of the Rashba SOC term, which in turn leads to the tuning and sign reversal of the THC.
Figure 5: Tuning the band topology with the Rashba SOC for the Néel type skyrmion crystal. (a) The Chern number of the lowest two subbands as a function of $\alpha_{R}$. (b, c, d) Evolution of the band structure with $\alpha_{R}$ changing through the crossover point $\alpha_{R}\approx-0.09$. The lowest two subbands touch and separate, and their Chern numbers change. (e, f) Berry curvatures for the lowest two subbands for $\alpha_{R}$ just above and below the crossover point, corresponding to the band structure of (b) and (d) respectively.
@@ III Topological Hall Effect with Rashba SOC
The THC in the skyrmion crystal without the Rashba SOC has been studied by earlier authors for several lattices. The results for the square lattice are shown in the Figure 3 both without and with the Rashba SOC term. When the Rashba SOC is absent, the Chern number for each of the subband is $\mathcal{C}_{n}$ of $-1$ as in the case of LLs arising from the uniform magnetic field [54, 55], except for the cluster of subbands around the van Hove singularity at $E=0$, which have the collective Chern number of $N-m$ as a group, where $N$ is the total number of subbands (16 here) and $m$ is the number of subbands in the cluster. If the Fermi level lies within a band gap, the THC is simply the negative of the sum of the Chern numbers of the lower bands in units of $e^{2}/h$. Due to the overlapping bands, these plateaus are not always visible.
@@
With the introduction of the Rashba SOC, the band dispersion is modified, and at an optimal value $\alpha_{R}^{0}$, the subbands become nearly dispersionless as discussed earlier. At this point, the THC quantization plateaus prominent, as seen in the case of quantum Hall effect (see Figure 3(b)).
@@
Not only the band dispersion is modified due to the presence of Rashba SOC, but the topology of the band structure is also altered, leading to a change in the Chern numbers. The Chern numbers still remain integers for the isolated individual subbands, but their magnitudes change as one varies the strength $\alpha_{R}$ of the Rashba SOC term. The results are shown for the $4\times 4$ square lattice in Figure 5 (a) for the lowest two subbands, where we see that the Chern numbers flip flop between the values ±$\pm 2$ for the range of $\alpha_{R}$ considered. Figure 5 (b-d) show the band structures as one passes through one of the crossover points, where the Chern number changes abruptly. The corresponding Berry curvatures are also shown in the Figure for $\alpha_{R}$ slightly below and above this crossover value. In essence, the change in the band topology occurs due to the changed magnetic field profile caused by the Rashba SOC term.
@@
Nanoribbon states. The varying nature of the band topology (Chern numbers) with changing $\alpha_{R}$ shows up in the edge states of the skyrmion crystal. We have computed the band structure in the nanoribbon geometry, where the ribbon is extended in the $x$ direction, while it has a finite width along the $y$ direction. The results are shown in Fig. 6 for three cases of $\alpha_{R}$. In particular, we focus on the lowest two bands and the edge states that occur in the midgap region between the two bands. The Chern numbers for the two lowest bulk subbands change from $C=-1$ to 0 to +1, as indicated in the top panel. The top panel also shows the edge states which are color coded red/green, corresponding to states localized on the bottom/top edge of the ribbon as seen from the middle panel of the figure. As the Chern numbers change from -1 to +1, the direction of propagation of the edge state is reversed, which is indicated by the long arrows in the middle panel of the Figure. For the case of $\alpha_{R}=-0.25$, Fig. 6(b), the lowest bulk bands are not topological in nature with zero Chern numbers, and the two edge states are degenerate in energy and constitute ordinary bound states.
@@
Band topology. The Chern numbers of the individual subbands change at the strength of the Rashba SOC $\alpha_{R}$ is varied, but always remaining an integer. The computed Chern numbers for the lowest two bands for four skyrmion types are shown in Fig. 7. The change in the Chern numbers is related to the change in the emergent magnetic field with $\alpha_{R}$. Except for the Néel skyrmion case, the Chern numbers are symmetric with respect to the change of the sign of $\alpha_{R}$, which can be understood in terms of the emergent magnetic fields, Eqs. (5-6), and the symmetry of the square lattice as discussed below.
@@
(1) For the case of Bloch skyrmion (with )helicity $\gamma=\pi/2$), the Chern numbers do not change with $\alpha_{R}$ as seen from Fig. 7 (c), since the $\cos\gamma=0$ factor in the magnetic field expression Eq. (5) removes any magnetic field dependence of $\alpha_{R}$.
@@
(2) For the antiskyrmions with any helicity, the Chern numbers are independent of the sign of $\alpha_{R}$ as seen from Figure 7 (b) and (d). This is because irrespective of the helicity for the antiskyrmions, the magnetic field, Eq. (6), is a $90^{\circ}$ rotation of each other when the sign of $\alpha_{R}$ is reversed due to the $\cos(2\alpha-\gamma)$ factor, while the square lattice has the same four-fold symmetry as well.
@@
(3) Unlike antiskyrmions, for the case of skyrmions, the Chern numbers for ±$\pm\alpha_{R}$ are not symmetric in general. This is because as Eq. (5) shows, the magnetic fields for ±$\pm\alpha_{R}$ are different and unrelated by any symmetry. An exception is the trivial case of $\cos\gamma=0$, where the Rashba contribution to the magnetic field Eq. (5)) is zero, which makes the Chern numbers entirely independent of $\alpha_{R}$, as seen from Fig. 7 (c).
@@
(4) Of the eight skyrmion types we have considered in this paper, Fig. 7 shows the Chern numbers only for the four cases. The other four cases in Figure 2 can be obtained from these results as well. Also, all properties for the Weyl SOC and Dresselhaus SOC including the subband Chern numbers can be expressed in terms of the results for the Rashba SOC as discussed later in the paper in Section V.

### v2

@@ Abstract
The topological Hall effect is the result of spin-asymmetric deflection of charge carriers flowing through a non-collinear spin system. Effective manipulation of the topological Hall conductivity (THC) in skyrmions is currently a vigorous area of research with an eye towards potential spintronics application. Here, we show that the band topology and the THC in a skyrmion crystal can be tuned by changing the strength of the Rashba spin-orbit coupling (SOC), which can be accomplished via a perpendicular electric field. This results in the change of the subband Chern numbers and a transition between ordinary insulator and Chern insulator as the Rashba SOC is varied. For partially filled subbands, the Rashba SOC can tune the THC and reverse its sign, so that the direction of the Hall current is flipped. The critical Rashba strength for this depends on the skyrmion type and the carrier density. We extend our analysis to the cases of Dresselhaus and Weyl SOC as well, and show that they can be directly mapped to the Rashba SOC case and therefore lead to similar results. Our work opens up the scope to go beyond the existing avenues for the control of charge transport in skyrmion crystals.
@@ I Introduction
Figure 2: The $4\times 4$ unit cell used in our calculation. Each site is a magnetic atom with its spin oriented according to the spin texture Eq. 4, with the origin being the center of the square. Together, the sixteen sites describe a single skyrmion, which is repeated via the lattice translation vectors $\mathbb{T_{1}}$ and $\mathbb{T_{2}}$ forming thereby the skyrmion crystal. In the figure, the central spin points into the page, the boundary spins point out of the page, while the rest of the spin directions are shown only schematically. The electron hops from atom to atom with a complex hopping amplitude due to the emergent magnetic field.
@@
In this work, we study the THE in the skyrmion crystal in the presence of three types of SOC, viz., the Rashba, Dresselhaus, or the Weyl type. The Rashba case is studied in some detail, while we show that the other two cases can be mapped to the Rashba case for the skyrmion types we have considered here. Quite remarkably, we find that not only can the magnitude of the topological Hall conductivity (THC) be significantly changed by varying the strength of the Rashba SOC via an applied electric field, but also that the sign of the Hall conductivity can be reversed by varying the Rashba SOC around a critical value. The band topology in the skyrmion crystal can also be changed by varying the strength of the SOC term, resulting in a change in the subband Chern numbers, which can lead to a flipping of the direction of the topological edge current.
@@ II Model Hamiltonian and the electron band structure
The first term is the nearest neighbor hopping term, where we have suppressed the spin index, with $c_{i}^{\dagger}\equiv(c_{i\uparrow}^{\dagger},c_{i\downarrow}^{\dagger})$ and $c_{i}\equiv(c_{i\uparrow},c_{i\downarrow})^{T}$, $i$ is the site index on the lattice, $\sigma$ is the spin index, $c_{i\sigma}^{\dagger}/c_{i\sigma}$ are the corresponding electron creation/annihilation operators, $\langle ij\rangle$ indicates distinct pairs of nearest neighbors, and H. c. indicates Hermitian conjugate of the last term only. Note that the sites represent individual magnetic atoms in the structure and several atoms taken together (16 here) form a single skyrmion, which is periodically repeated to make the skyrmion crystal (see Figure 2). The second term is the exchange interaction between the site-fixed spins $\bm{S}_{i}$, commonly taken as classical spins, with the skyrmion texture and $\bm{\sigma}$ is the Pauli matrix describing the electron spin. The third term is the lattice version of the Rashba SOC[46, 47], which breaks the $z\rightarrow-z$ symmetry. In this term, $\alpha_{R}$ is the Rashba interaction strength and $\mathbb{r}_{ij}=\mathbb{r}_{j}-\mathbb{r}_{i}$ is the vector distance between the nearest-neighbor sites. It is easily shown that for the free-electron case (small $k$ limit of the lattice case), it yields the well-known form of the Rashba Hamiltonian $^{\cal H}_{R}^{k}=\alpha_{R}^{\prime}(\bm{\sigma}\times\bm{k})\cdot\hat{z}$, with a scaled coefficient $\alpha_{R}^{\prime}=2a^{2}\alpha_{R}$, where $a$ is the lattice constant [48]. The Rashba SOC can arise due to the symmetry breaking by an overlayer or by applying an external electric field.
For the free electrons in the presence of an electric field, the strength of the Rashba term is given by $\alpha_{R}^{\prime}=\hbar^{2}E_{z}/2m_{e}^{2}c^{2}$. This expression however severely underestimates the magnitude of the Rashba SOC, if we take $E_{z}$ to be the surface electric field. The reason for this discrepancy[49] is that the relativistic effects originate largely from the nuclear region, and the role of the surface electric field is to modify the electron density near the nucleus by mixing the orbitals due to broken symmetry. This is a much larger effect than simply computing a free-electron-like expression due to the surface electric field.
The exchange interaction is often taken as $J\rightarrow\infty$, which is reasonable considering that for typical solids $J\sim 2-3$ eV, while the typical hopping is a fraction of an eV. Also, the essential physics discussed here is not changed by taking this limit. However, there are no additional complications if a general $J$ is used. As an example, the band structure and THC for finite $J$ are shown in the Supplementary Material [48]. In infinite $J$ limit, the problem reduces to that of the spinless fermions, since the electron state with spin anti-aligned with the lattice spin is forbidden. The spinless fermion operators are described by
@@
where $(\theta,\phi)$ are the polar and azimuthal angles in the spherical polar coordinates and $(r,\alpha)$ is the 2D position of the lattice spin in polar coordinates with the origin coinciding with the skyrmion centre. Here, $\lambda$ is the skyrmion radius, $m$ is the winding number or the vorticity, and $\gamma$ is the helicity of the skyrmion. For a single-turned skyrmion and antiskyrmion, $m$ takes the value of +1+1 and 1-1, respectively. Depending on the helicity, skyrmions can be classified as Néel ($\gamma=0\ {\rm or}\ \pi$) and Bloch types ($\gamma=\pm\pi/2$) as illustrated in Figure 3.
@@
We consider a square lattice of localized spins with a magnetic unit cell of $N\times N$ sites hosting a skyrmion with the spin texture Eq. 4. Thus the skyrmions form a square lattice with lattice constant $Na$, where $a=1$ is the distance between the neighboring spins. We took $N=4$ with the skyrmion radius $\lambda=2a$ in our calculation, so that there are $N^{2}=16$ sites in the magnetic unit cell of the skyrmion crystal. The unit cell of the skyrmion crystal is shown in Figure 2. Here, we have used a smaller ($4\times 4$) magnetic cell for illustration, but the physics remains unchanged if a larger cell is used, since the total flux through the cell remains the same. Some results for the $10\times 10$ are shown in the Supplementary Materials [48].
@@
The SOC modifies the emergent magnetic field of the skyrmion, which we discuss below. In addition, it modifies the magnitude of the effective hopping [48]. Both these effects modify the electronic band structure as well as the topological properties.
Emergent magnetic field. Here, we compute the emergent magnetic field seen by the electron for an isolated skyrmion in the presence of the Rashba SOC. The results will be needed in the interpretation of the THC, particularly the quantum Hall plateaus, as the SOC strength is varied.
@@
In the lattice models with a magnetic field, hopping integrals between lattice sites acquire a complex phase factor $t\times\exp\big{[}\frac{-ie}{\hbar}\int_{a}^{b}\mathbb{A}.d\mathbb{r}\big{]}$ via the Peierls substitution, where $t$ is the hopping without the magnetic field from site $a$ to $b$, $\mathbb{A}$ is the vector potential, e<0-e<0 is the charge of the electron, and the integral is taken over the path joining the two sites. To compute the magnetic field, we take an infinitesimal square loop around the point $\mathbb{r}$ and compute the flux through it using the Peierls phase factors accumulated by the electron as it travels around the square loop, following the Hamiltonian Eq. (3). After straightforward but tedious algebra [48], we find that the magnetic field, which is normal to the skyrmion plane, is given by
@@
$B_{SK}^{R}(r)=\begin{cases}\frac{\Phi_{0}}{2\pi}\times\big{[}\frac{\pi}{2\lambda r}\sin\frac{\pi r}{\lambda}-\frac{\alpha_{R}}{t}(\frac{\pi}{\lambda}\cos{\frac{\pi r}{\lambda}}+\frac{1}{r}\sin\frac{\pi r}{\lambda})\cos\gamma\big{]}&\text{for}~r<\lambda\\ 0&\text{for}~r>\lambda,\end{cases}$ (5)
where $\Phi_{0}=h/e$ is the flux quantum. The superscript in $B_{SK}^{R}$ refers to the Rashba SOC case. The first term is due to the skyrmion alone, and it is independent of the helicity $\gamma$, but depends on the vorticity ($m=+1$ here), while the second term is the modification of the magnetic field due to the Rashba SOC. The $B_{SK}^{R}$ has a discontinuity at the boundary which arises due to the discontinuity in the derivative of the spin texture Eq. 4. This can be understood from the magnetic field expression of Eq. 20 discussed later.
For the antiskyrmion, the magnetic field due to Rashba SOC is
@@
$B_{AS}^{R}(r,\alpha)=\begin{cases}\frac{\Phi_{0}}{2\pi}\times\big{[}-\frac{\pi}{2\lambda r}\sin\frac{\pi r}{\lambda}-\frac{\alpha_{R}}{t}(\frac{\pi}{\lambda}\cos{\frac{\pi r}{\lambda}}-\frac{1}{r}\sin\frac{\pi r}{\lambda})\cos({2\alpha}-\gamma)\big{]}&\text{for}~r<\lambda\\ 0&\text{for}~r>\lambda,\end{cases}$ (6)
@@
Note that here the magnetic field is no longer circularly symmetric. It can be shown by direct integration that the total flux passing through a single syrmion/ antiskyrmion is $\pm\Phi_{0}$, irrespective of the strength of the Rashba term. In other words, the Rashba term does not change the total flux through the skyrmion, rather, it redistributes the flux within the skyrmion region, so that the total flux remains unchanged. The same is also true for the Weyl and Dresselhaus SOC as well.
@@
Figure 4: Band structure and THC in the skyrmion crystal with the Néel skyrmion. (a) Band structure without (red lines) and with (blue lines) Rashba SOC. For the latter case, only the lowest four bands are shown. (b) THC as a function of the Fermi energy $\varepsilon_{F}$, indicating prominent Hall plateaus for $\alpha_{R}=0.5t/a$, chosen so as to produce a maximally uniform magnetic field (see Figure 5). Here and throughout the paper, the Hund’s coupling is taken to be $J\rightarrow\infty$ as discussed in the text.
@@
We note that the electron states in the skyrmion crystal are in many ways similar to the electron states in the integer quantum Hall effect (IQHE), in the sense that while the electrons are subjected to a uniform magnetic field in the IQHE, for the skyrmion crystal, the emergent magnetic field is non-uniform and varies spatially. The band structures as well as the topological properties for the two cases are similar (a specific example is given in the Figure 1 of the Supplementary Materials [48]).
@@
To obtain this optimal value $\alpha_{R}^{0}$ that produces the maximally uniform magnetic field, we have minimized the smoothness integral $\int_{0}^{\lambda}2\pi r\left(dB_{SK}(r)/dr\right)^{2}dr$ over the skyrmion radius using the total magnetic field Eq. (5). The result is:
@@
where $c=c_{1}-c_{2}+\ln\ (2\pi)$, $c_{1}\approx 0.577$ is the Euler-Mascheroni constant, and $c_{2}=-\int_{2\pi}^{\infty}dx\ \cos x/x\approx-0.022$ is the cosine integral. From an inspection of Eq. (5), it is obvious that the magnetic field can never be uniform for all $r$, but for the optimal Rashba strength $\alpha_{R}^{0}$, we find that the rms deviation of $B(r)$ from the mean is as small as $\approx 8\%$. Figure 5 shows the magnetic field for the optimal value $\alpha_{R}^{0}$.
@@
The band structure for the $4\times 4$ lattice with the Rashba strength $\alpha_{R}^{0}$ is shown in Figure 4 (a) (in blue), which is very similar to the bands with a uniform magnetic field [48]. As discussed in the next Section, we show that the topology of the bands becomes modified in the presence of the Rashba SOC term, which in turn leads to the tuning and sign reversal of the THC.
Figure 6: Tuning the band topology with the Rashba SOC for the Néel type skyrmion crystal. (a) The Chern number of the lowest two subbands as a function of $\alpha_{R}$. (b, c, d) Evolution of the band structure with $\alpha_{R}$ changing through the crossover point $\alpha_{R}\approx-0.09$. The lowest two subbands touch and separate, and their Chern numbers change as indicated in (b) and (d), where we have shown the Chern numbers of the two lowest subbands. (e, f) Berry curvatures for the lowest two subbands for $\alpha_{R}$ just above and below the crossover point, corresponding to the band structure of (b) and (d) respectively.
@@ III Topological Hall Effect with Rashba SOC
The THC in the skyrmion crystal without the Rashba SOC has been studied by earlier authors for several lattices. The results for the square lattice are shown in the Figure 4 both without and with the Rashba SOC term. When the Rashba SOC is absent, the Chern number for each of the subband is $\mathcal{C}_{n}$ of 1-1 as in the case of LLs arising from the uniform magnetic field [52, 53], except for the cluster of subbands around the van Hove singularity at $E=0$, which have the collective Chern number of NmN-m as a group, where $N$ is the total number of subbands (16 here) and $m$ is the number of subbands in the cluster. If the Fermi level lies within a band gap, the THC is simply the negative of the sum of the Chern numbers of the lower bands in units of $e^{2}/h$. Due to the overlapping bands, these plateaus are not always visible.
@@
With the introduction of the Rashba SOC, the band dispersion is modified, and at an optimal value $\alpha_{R}^{0}$, the subbands become nearly dispersionless as discussed earlier. At this point, the THC quantization plateaus prominent, as seen in the case of quantum Hall effect (see Figure 4(b)). Due to the quantization in the Hall conductivity, it is sometimes referred to as quantized topological Hall effect (QTHE)[52].
@@
Not only the band dispersion is modified due to the presence of Rashba SOC, but the topology of the band structure is also altered, leading to a change in the Chern numbers. The Chern numbers still remain integers for the isolated individual subbands, but their magnitudes change as one varies the strength $\alpha_{R}$ of the Rashba SOC term. The results are shown for the $4\times 4$ square lattice in Figure 6 (a) for the lowest two subbands, where we see that the Chern numbers flip flop between the values $\pm 2$ for the range of $\alpha_{R}$ considered. Figure 6 (b-d) show the band structures as one passes through one of the crossover points, where the Chern number changes abruptly. The corresponding Berry curvatures are also shown in the Figure for $\alpha_{R}$ slightly below and above this crossover value. In essence, the change in the band topology occurs due to the changed magnetic field profile caused by the Rashba SOC term.
Finally, we note that when the topological character (Chern number) of a band changes as a function of any variable ($\alpha_{R}$ here), the system must go through a metallic state, as seen from Figure 6 (b), (c), and (d).
@@
Nanoribbon states. The varying nature of the band topology (Chern numbers) with changing $\alpha_{R}$ shows up in the edge states of the skyrmion crystal. We have computed the band structure in the nanoribbon geometry, where the ribbon is extended in the $x$ direction, while it has a finite width along the $y$ direction. The results are shown in Figure 7 for three cases of $\alpha_{R}$. In particular, we focus on the lowest two bands and the edge states that occur in the midgap region between the two bands. The Chern numbers for the two lowest bulk subbands change from C=1C=-1 to 0 to +1, as indicated in the top panel. The top panel also shows the edge states which are color coded red/green, corresponding to states localized on the bottom/top edge of the ribbon as seen from the middle panel of the figure. As the Chern numbers change from -1 to +1, the direction of propagation of the edge state is reversed, which is indicated by the long arrows in the middle panel of the Figure. For the case of $\alpha_{R}=-0.25$, Figure 7(b), the lowest bulk bands are not topological in nature with zero Chern numbers, and the two edge states are degenerate in energy and constitute ordinary bound states.
@@
Band topology. The Chern numbers of the individual subbands change as the strength of the Rashba SOC $\alpha_{R}$ is varied, but always remaining an integer. The computed Chern numbers for the lowest two bands for four skyrmion types are shown in Figure 8. The change in the Chern numbers is related to the change in the emergent magnetic field with $\alpha_{R}$. Except for the Néel skyrmion case, the Chern numbers are symmetric with respect to the change of the sign of $\alpha_{R}$, which can be understood in terms of the emergent magnetic fields, Eqs. (5-6), and the symmetry of the square lattice as discussed below.
@@
(1) For the case of Bloch skyrmion (with helicity $\gamma=\pi/2$), the Chern numbers do not change with $\alpha_{R}$ as seen from Figure 8 (c), since the $\cos\gamma=0$ factor in the magnetic field expression Eq. (5) removes any magnetic field dependence of $\alpha_{R}$.
@@
(2) For the antiskyrmions with any helicity, the Chern numbers are independent of the sign of $\alpha_{R}$ as seen from Figure 8 (b) and (d). This is because irrespective of the helicity for the antiskyrmions, the magnetic field, Eq. (6), is a $90^{\circ}$ rotation of itself when the sign of $\alpha_{R}$ is reversed due to the $\cos(2\alpha-\gamma)$ factor, while the square lattice has the same four-fold symmetry as well.
@@
(3) Unlike antiskyrmions, for the case of skyrmions, the Chern numbers for $\pm\alpha_{R}$ are not symmetric in general. This is because as Eq. (5) shows, the magnetic fields for $\pm\alpha_{R}$ are different and unrelated by any symmetry. An exception is the trivial case of $\cos\gamma=0$, where the Rashba contribution to the magnetic field Eq. (5)) is zero, which makes the Chern numbers entirely independent of $\alpha_{R}$, as seen from Figure 8 (c).
@@
(4) Of the eight skyrmion types we have considered in this paper, Figure 8 shows the Chern numbers only for the four cases. The other four cases in Figure 3 can be obtained from these results as well. Also, all properties for the Weyl SOC and Dresselhaus SOC including the subband Chern numbers can be expressed in terms of the results for the Rashba SOC as discussed later in the paper in Section V.